5. A fair die is rolled until a 6 occurs. Compute the probability that (a) 10 rolls are needed.

(b) an odd number of rolls is needed.Β 


1
Expert's answer
2021-03-08T17:14:06-0500

a) The probability of getting 6 in the first roll P(1) is "\frac{1} {6}"


The probability of getting 6 in the second roll P(2) is "\frac{5}{6}×\frac{1} {6}"


Thus the probability of P(10);

"\frac{5}{6} ×\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{1 }{6}"


="\frac{4}{125}"


b)an odd number P(1) P(3)P(5)


"\frac{5}{6}×\frac{5}{6}×\frac{5}{6}×\frac{1 }{6}"


="\frac{125}{1296}"